A room is divided by a rectangle partition 6 feet high and 8 feet long. How many square tiles, each of which covers a 6 inch by 6 inch area, are required to cover one side of this partition? I don't know how to even start!
Answer by checkley75(3666)

THE AREA TO BE COVERED IS 6*8=48 FEET SQUARED.
EACH TILE IS 6*6=36 SQUARE INCHES.
A SQUARE FOOT=12*12=144 SQUARE INCHES.
THEREFORE THERE ARE 144/36= TILES PER SQUARE FOOT.4
THEREFORE THERE ARE 4*48=192 TILES NEEDED TO COVER THE RECTANGULAR PARTITION.